### v1

# Spectrality of Prime Size Tiles

Abstract
We prove that if a tile in $\mathbb{Z}^{d}$ has prime size $p$, then it must be spectral. The proof is by contradiction, it is simply shown that the tiling complement of such a tile can not annihilate all $p$-subgroups. In addition, with a simple transformation we prove that any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
Keywords: Fuglede conjecture; tiling sets; spectral sets.
2020 MSC: 42A99; 05B45
1 Introduction
Let $G$ be a finitely generated Abelian group, a set $A$ is tiling (or is a tile) in $G$ if there exists some $B$ such that $\{A+b\}_{b\in B}$ forms a partition of $G$, and is spectral if there exists a set of group characters of $G$ that forms an orthogonal basis on $L^{2}(A)$. In such cases $B$ (resp. $S$) is said to be a tiling complement (resp. spectrum) of $A$, while $(A,B)$ (resp. $(A,S)$) is called a tiling pair (resp. spectral pair) in $G$, and we often write $G=A\oplus B$.
Fuglede conjectured that being spectral and being tiling are actually equivalent [6], this has been studied extensively and disproved in general when the dimension is at least three [5, 10, 11, 17, 20] but remains open for one dimensional and two dimensional cases. Researches over the Fuglede conjecture in finite Abelian groups as well as in $p$-adic fields have been active in the past years, see e.g., [3, 4, 9, 12, 13, 15, 16, 21] (the full list is of course too long to be enumerated here, this is only an attemp to mention some of the results in one dimensional or two dimensional cases, see perhaps references therein for more on the history and development of this problem).
It is well known that a tiling set of prime size in $\mathbb{Z}^{d}$ must have at least one periodic tiling complement [19] (it may be worthing mentioning that being periodic in $d$ dimensions means having $d$ linearly independent periods, while having periods in one direction is called weakly periodic instead, readers may also find the periodic tiling conjecture and its positive and negative cases in e.g., [1, 7, 8, 14] relevant). The aim of this short note is to establish the following statements (Theorem 1 and 2 in the last section):
Theorem.
Let $p$ be a prime number and $A\subset\mathbb{Z}^{d}$ with $|A|=p$, if $A$ tiles $\mathbb{Z}^{d}$, then it is also spectral.
Theorem.
Any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is acutually already included in [17, Corollary 2.4].
The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$. The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, while the existence of such a set being included in the zero set of $A$ is shown by contradiction. Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$, thus leading to a contradiction if $A$ does not admit such a spectrum.
The proof of the second theorem then reduces to proving any $p$ points in general linear positions are tiling, for which we first apply a group isomorphism to map them to the origin and $(1,0\ldots,0),(0,2,0,\ldots,0),\ldots,(0,\ldots,0,p-1)$, then we construct a tiling complement for such a set explicitly.
2 Preliminaries
Let $A$ be a multiset (i.e., a set that allows repetitive elements) on $\mathbb{Z}_{n}^{d}$ (i.e., each element of $A$ is a member of $\mathbb{Z}_{n}^{d}$), the multiplicity of an element $a\in A$ (i.e., the number of copies of $a$ in $A$) will be denoted by $\operatorname{mul}(a)$, the characteristic function on such a multiset $A$ will be defined and written as
$\mathbf{1}_{A}(x)=\begin{cases}\operatorname{mul}(x)&x\in A,\\ 0&x\notin A.\end{cases}$
The Fourier transform of $\mathbf{1}_{A}(x)$ on $\mathbb{Z}_{n}^{d}$ is
$\widehat{\mathbf{1}_{A}}(\xi)=\sum_{a\in A}\operatorname{mul}(a)\cdot e^{2\pi i\langle a,\xi\rangle/n},$
where $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. It is straigtforward to verify that
[…]
where the difference set $\Delta S$ of a usual set (i.e., a set with distinct elements only) $S$ and the zero set $Z(f)$ of a function $f$ is defined respectively as
$\Delta S=\{s-s^{\prime}:s,s^{\prime}\in S,\;s\neq s^{\prime}\},\qquad Z(f)=\{x:f(x)=0\}.$
One may also verify the Poisson summation formula below:
$\widehat{\mathbf{1}_{H}}=|H|\cdot\mathbf{1}_{H^{\perp}},$ (3)
where $H\lhd\mathbb{Z}_{n}^{d}$ is a subgroup and
$H^{\perp}=\{x\in\mathbb{Z}_{n}^{d}:\langle x,h\rangle=0,\forall h\in H\},$
is its orthogonal group in $\mathbb{Z}_{n}^{d}$. That(3) holds for cyclic subgroups is quite trivial, while non-cyclic subgroups can be viewed as products of cyclic ones, its orthogonal set is then the intersection over orthogonal sets of its cyclic factors.
Let the support of a function $f$ in $\mathbb{Z}_{n}^{d}$ be defined as
$\operatorname{supp}(f)=\mathbb{Z}_{n}^{d}\setminus Z(f),$
then the uncertainty principle on finite Abelian groups [2, 18] asserts that
$|\operatorname{supp}(f)|\cdot|\operatorname{supp}(\widehat{f})|\geq|\mathbb{Z}_{n}^{d}|,$ (4)
holds for any non-zero function $f$ defined on $\mathbb{Z}_{n}^{d}$.
Lemma 1.
Let $A$ be a multiset on $\mathbb{Z}_{n}^{d}$, if $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{A})$, then $\mathbf{1}_{A}$ must be a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, i.e., $\mathbf{1}_{A}=m\mathbf{1}_{\mathbb{Z}_{n}^{d}}$ for some $m\in\mathbb{N}$.
Proof.
Assume the contrary that $\mathbf{1}_{A}$ is not a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, then let $m$ be the smallest multipicity of elements in $A$, i.e.,
$m=\min\{\operatorname{mul}(a):a\in A\},$
and consider
$f=\mathbf{1}_{A}-m\mathbf{1}_{\mathbb{Z}_{n}^{d}}.$
With this construction we would have $f(a)=0$ for all $a$ whose multiplicity in $A$ is $m$ and $f$ is not identically 0 (since $\mathbf{1}_{A}$ is assumed not to be a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$). Therefore we get
$0<|\operatorname{supp}(f)|<|\mathbb{Z}_{n}^{d}|,$ (5)
On the other hand, it is straightforward to check that
$Z(\widehat{\mathbf{1}_{\mathbb{Z}_{n}^{d}}})=\mathbb{Z}_{n}^{d}\setminus\{0\}=\Delta\mathbb{Z}_{n}^{d},$
i.e., $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{\mathbb{Z}_{n}^{d}})$, thus combined with the assumption that $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{A})$ we get
[…]
i.e.,
$\operatorname{supp}(\widehat{f})=\{0\}.$
Together with (5) this leads to
$|\operatorname{supp}(f)|\cdot|\operatorname{supp}(\widehat{f})|=|\operatorname{supp}(f)|\cdot 1=|\operatorname{supp}(f)|<|\mathbb{Z}_{n}^{d}|,$
which contradicts the uncertainty principle in (4). ∎
For any $x=(x_{1},\ldots,x_{d})\in\mathbb{Z}^{d}$, set
$\pi_{n}(x)=(x_{1}\bmod n,\;\ldots,\;x_{d}\bmod n),$
so that $\pi_{n}$ is the projection from $\mathbb{Z}^{d}$ to $\mathbb{Z}_{n}^{d}$ (as a map between sets). Given a set $A\subseteq\mathbb{Z}^{d}$, $\pi_{n}(A)$ will denote the multiset obtained by applying $\pi_{n}$ on every element of $A$. Clearly $\pi_{n}(A)$ will only be a usual set if (and only if) the map $A\mapsto\pi_{n}(A)$ is injective.
Lemma 2 ([19]).
Let $p$ be a prime number and $A\subset\mathbb{Z}^{d}$ with $|A|=p$, if $A\oplus B=\mathbb{Z}^{d}$, then there is some $n\in\mathbb{N}$ and $B^{\prime}\in\mathbb{Z}_{n}^{d}$ such that $A\mapsto\pi_{n}(A)$ is injective and $\pi_{n}(A)\oplus B^{\prime}=\mathbb{Z}_{n}^{d}$.
The way that Lemma 2 is stated here is slightly different from its original version in [19, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.
Lemma 3.
Let $n=p^{k}m$ with $p$ being a prime number that does not divide $m$, and set $A^{\prime}=\pi_{p^{k}}(A)$. Let both $f,g$ be characteristic functions on $A^{\prime}$ but with the domain of $f$ being the entire $\mathbb{Z}_{n}^{d}$ while the domain of $g$ being $\mathbb{Z}_{p^{k}}^{d}$ only. If $x\in\mathbb{Z}_{n}^{d}$ is an element of order $p^{t}$ for some $1\leq t\leq k$ , then we have
$\widehat{\mathbf{1}_{A}}(x)=\widehat{f}(x)=\widehat{g}(x^{\prime}),$
where $x^{\prime}\in\mathbb{Z}_{p^{k}}^{d}$ is the element that satisfies $x=mx^{\prime}$, and the last Fourier transform $\hat{g}$ is performed in $\mathbb{Z}_{p^{k}}^{d}$ only (the first two Fourier transforms $\widehat{\mathbf{1}_{A}}$ and $\widehat{f}$ take place in $\mathbb{Z}_{n}^{d}$).
Proof.
Every $a=(a_{1},\ldots,a_{d})\in A$ can be written as
$a=p^{k}q_{a}+r_{a},$
where $q_{a}$, $r_{a}$ respectively collect quotients and remainders of $a_{1},\ldots,a_{d}$ divided by $p^{k}$. In particular we have
$A^{\prime}=\pi_{p^{k}}(A)=\{r_{a}:a\in A\}.$
The support set of $f,g$ coincide, but they are defined on different subgroups, thus we have
$\widehat{f}(\xi)=\sum_{a\in A}e^{2\pi i\langle r_{a},\xi\rangle/n},\qquad\widehat{g}(\xi)=\sum_{a\in A}e^{2\pi i\langle r_{a},\xi\rangle/p^{k}}$
in which Fourier transforms in $\mathbb{Z}_{n}^{d}$ and in $\mathbb{Z}_{p^{k}}^{d}$ are applied respectively.
Then direct computation shows that
$\widehat{\mathbf{1}_{A}}(x)$ $=\sum_{a\in A}e^{2\pi i\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\langle p^{k}q_{a}+r_{a},mx^{\prime}\rangle/n}=\sum_{a\in A}e^{2\pi i\langle r_{a},mx^{\prime}\rangle/n},$
$=\sum_{a\in A}e^{2\pi i\langle r_{a},x\rangle/n}\left(=\widehat{f}(x)\right),$
$=\sum_{a\in A}e^{2\pi i\langle r_{a},x^{\prime}\rangle/p^{k}}\left(=\widehat{g}(x^{\prime})\right),$
which is the desired result. ∎
Given a finite Abelian group $G$, let $\sim$ be the equivalence relation on $G$ so that $g\sim g^{\prime}$ if $g$ and $g^{\prime}$ generate the same cyclic subgroup in $G$. If $E$ is an equivalence class under $\sim$, then we define the order of $E$, denoted by $\operatorname{ord}(E)$, to be the order of any element (as a member of the group $G$) in $E$ (by definition all elements in the same class will have the same order).
Lemma 4.
Let $E$ be an equaivlence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $A$ annihilates one element of $E$, then it annihilates entire $E$.
Proof.
Set $m=\operatorname{ord}(E)$ and $r=n/m$. Let $x$ be an arbitrary element of $E$, then $x=rx^{\prime}$ for some $x^{\prime}\in\mathbb{Z}_{m}^{d}$, consequently
$\widehat{1_{A}}(x)=\sum_{a\in A}e^{2\pi i\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\langle a,x^{\prime}\rangle/m}.$
We shall view the above expression as the polynomial
$P(z)=\sum_{a\in A}z^{\langle a,x^{\prime}\rangle},$
evaluated at the $m$-th root of unity $\omega=e^{2\pi i/m}$, i.e.,
$\widehat{1_{A}}(x)=P(\omega).$
[…]
$\widehat{1_{A}}(y)=\sum_{a\in A}e^{2\pi i\langle a,\lambda x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x^{\prime}\rangle/m}=P(\omega^{\lambda})=0,$
where the last equality holds since $\lambda$ is coprime to $m$, and so that $\omega^{\lambda}$ is also a root of $\Phi_{m}(z)$. This finishes the proof since $x,y$ are chosen arbitrarily from $E$. ∎
A set $E^{\prime}$ will be called a derived set of $E$ if
$E^{\prime}\subseteq E\cup\{0\},\qquad\Delta E^{\prime}\subseteq E.$
Lemma 5.
Let $E$ be an equaivlence class under $\sim$ in some finite Abelian group. If $p$ is the smallest prime divisor of $\operatorname{ord}(E)$, then $E$ admits a derived set of size $p$.
Proof.
Let $h$ be an arbitrary element of $E$, then we claim that
$E^{\prime}=\{0,h,2h,\ldots,(p-1)h\},$
is a desired derived set. Indeed, that $|E^{\prime}|=p$ is obvious, to see $\Delta E^{\prime}\subseteq E$, set $n=\operatorname{ord}(E)$ and simply consider the group isomorphism from the cyclic group generated by $h$ to $\mathbb{Z}_{n}$ given by $\varphi:h\mapsto 1$. Let $F$ be the equivalence class formed by taking all generators of $\mathbb{Z}_{n}$, then $F$ consists of all numbers that are coprime to $n$ in $\mathbb{Z}_{n}$ and by this construction we have $\varphi(E)=F$. Now set $F^{\prime}=\{0,1,\ldots,p-1\}$, then $\Delta F^{\prime}=\{\pm 1,\ldots,\pm(p-1)\}$, clearly every member of $\Delta F^{\prime}$ is coprime to $n$ since $p$ is the smallest divisor of $n$. Therefore $\Delta F^{\prime}\subseteq F$, apply $\varphi^{-1}$ at both sides we get $\Delta E^{\prime}\subseteq E$. ∎
In particular, Lemma 5 holds for $\operatorname{ord}(E)$ being a power of $p$, which is how it will be applied later in this article.
3 Main results
Theorem 1.
Let $p$ be a prime number and $A\subset\mathbb{Z}^{d}$ with $|A|=p$, if $A$ tiles $\mathbb{Z}^{d}$, then it is also spectral.
Proof.
Let $n$ and $B^{\prime}$ be as asserted in Lemma 2, and set $A^{\prime}=\pi_{n}(A)$, so that we have $A^{\prime}\oplus B^{\prime}=\mathbb{Z}_{n}^{d}$. It suffices to show that $A^{\prime}$ is spectral in $\mathbb{Z}_{n}^{d}$.
$A^{\prime}$ being a tile in $\mathbb{Z}_{n}^{d}$ implies that $p$ divides $n^{d}$. Since $p$ is prime this also implies that $p$ divides $n$. Let $m$ be the number so that $n=p^{k}m$ and $\gcd(p,m)=1$.
[…]
By Lemma 5, each of these equivalence classes has a derived set of size $p$, thus if any of these classes is in $Z(\widehat{\mathbf{1}_{A^{\prime}}})$, then the corresponding derived set would be a spectrum of $A^{\prime}$ and we are done.
To see this must be the case, assume the contrary that all equivalence classes whose orders are powers of $p$ are disjoint with $Z(\widehat{\mathbf{1}_{A^{\prime}}})$, then by (2) they must be annihilated by $\widehat{\mathbf{1}_{B^{\prime}}}$. By Lemma 3 this further implies
[…]
Indeed, every element $x^{\prime}\in\Delta\mathbb{Z}_{p^{k}}^{d}$ can be lifted to an element of the same order in $\mathbb{Z}_{n}^{d}$ through the map $x^{\prime}\mapsto mx^{\prime}$, and Lemma 3 asserts that
$\widehat{\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}}(x^{\prime})=\widehat{\mathbf{1}_{B^{\prime}}}(mx^{\prime}).$ (7)
The order of $x^{\prime}$ in $\mathbb{Z}_{p^{k}}^{d}$ is obviously a power of $p$, thus the order of $mx^{\prime}$ in $\mathbb{Z}_{n}^{d}$ is also a power of $p$ (since it equals the order of $x^{\prime}$ in $\mathbb{Z}_{p^{k}}^{d}$), which means by assumption the right-hand side of (7) must vanish. Repeating the argument on every member of $\Delta\mathbb{Z}_{p^{k}}^{d}$ leads to (6).
Now as $\pi_{p^{k}}(B^{\prime})$ is a multiset on $\mathbb{Z}_{p^{k}}^{d}$, Lemma 1 indicates that $\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}$ is just a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{p^{k}}^{d}}$, which means $|Z_{p^{k}}^{d}|$ divides $|\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}|$. Since $|\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}|=|B^{\prime}|$, this further indicates that $p^{kd}$ divides $|B^{\prime}|$, consequently we would have $p^{kd}m^{d}=|\mathbb{Z}_{n}^{d}|=|A|\cdot|B|$ is divisible by $p^{kd+1}$, which is a contradiction. ∎
Lemma 6.
Given any $n$ linearly indpenedent vectors $v_{1},\ldots,v_{n}$ on $\mathbb{Z}^{n}$, there exists a group automorphism of the additive group $\mathbb{Z}^{n}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,1,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,1)$. In particular, if $p$ is a prime number and n=p1n=p-1, then there exists a group automorphism of the additive group $\mathbb{Z}^{p-1}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$.
Proof.
Let $V$ be the matrix whose columns are $v_{1},\ldots,v_{n}$, and $D$ the determinant of $V$. Clearly $D\neq 0$ because of linear independence.
Let $V^{\prime}$ be the adjugate matrix of $V$, i.e., if $U_{i,j}$ is the determinant of the matrix obtained by deleting the $i$-th row and the $j$-th column of $V$, then $V^{\prime}$ is the transpose of the matrix whose $i,j$-th entry is $(-1)^{i+j}U_{i,j}$. Then $V^{\prime}V=DI$, and thus $V^{\prime}$ is a group isomorphism from $\mathbb{Z}^{n}$ to $D\mathbb{Z}^{n}$ that maps $v_{1},\ldots,v_{p-1}$ to $(D,0,\ldots,0)$, $(0,D,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,D)$. The desired automorphism can then be constructed by composing the following two isomorphisms:
$\mathbb{Z}^{n}\xrightarrow{\qquad\qquad V^{\prime}\qquad\qquad}D\mathbb{Z}^{n}\xrightarrow{\qquad x\mapsto x/D\qquad}\mathbb{Z}^{n},$
If n=p1n=p-1, then $\operatorname{diag}(1,2,\ldots,{p-1})$ is also a group autophism on $\mathbb{Z}^{p-1}$, thus the desired automorphism is constructed by a further composition:
$\mathbb{Z}^{p-1}\xrightarrow{\qquad\qquad\qquad V^{\prime}\qquad\qquad\qquad}D\mathbb{Z}^{p-1}\xrightarrow{\qquad\qquad x\mapsto x/D\qquad\qquad}\mathbb{Z}^{p-1}\xrightarrow{\qquad\operatorname{diag}(1,2,\ldots,{p-1})\qquad}\mathbb{Z}^{p-1}.$
Lemma 7.
The set
$A=\{(0,0,\ldots,0),(1,0,\ldots,0),(0,2,0,\ldots,0),\ldots,(0,\ldots,0,p-1)\},$
is both tiling and spectral in $\mathbb{Z}_{p}^{p-1}$.
Proof.
Let $x=(1,\ldots,1)$, and $\omega=e^{2\pi i/p}$, then it is easy to verify that
$\widehat{\mathbf{1}_{A}}(x)=1+\omega+\ldots+\omega^{p-1}=0.$
Hence the cyclic group generated by $x$ (denoted by $\langle x\rangle$) is a spectrum of $A$. To see that $A$ is tiling, we will construct a set $B$ so that $|A|\cdot|B|=|\mathbb{Z}_{p}^{p-1}|$ and
$Z(\widehat{\mathbf{1}_{B}})=\mathbb{Z}_{p}^{p-1}\setminus\langle x\rangle,$ (8)
then it will follow from (2) that $(A,B)$ forms a tiling pair.
Since $p$ is a prime number, the (p1)(p-1)-st roots of unity in the finite field $\mathbb{F}_{p}$ is precisely $1,\ldots,p-1$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and its columns, denoted by $u_{1},\ldots,u_{p-1}$ is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$). In particular, since $x$ is a column of it we may without loss of generality take $u_{1}=x$.
Let $B$ be the span (over $\mathbb{F}_{p}$) of $u_{2},\ldots,u_{p-1}$, then $|B|=p^{p-2}$ is of the correct size, and we simply notice that $B$ and $\langle x\rangle$ are orthogonal groups of each other, thus (8) follows immediately from the Poisson summation formula (3). ∎
Theorem 2.
Any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
Proof.
In light of Theorem 1, it suffices to show that these points are tiling in $\mathbb{Z}^{d}$.
Without loss of generality we may assume that one of the point is the origin, and the other p1p-1 points are $v_{1},\ldots,v_{p-1}$. By Lemma 6, we can find a group automorphism of $\mathbb{Z}^{d}$ that maps the original $p$ points $0,v_{1},\ldots,v_{p-1}$ to
$A=\{0,e_{1},2e_{2},\ldots,(p-1)e_{p-1}\},$
where $e_{i}$ is the $i$-th standard Euclidean basis. Since it is a group automorphism we assert that $A$ is tiling if and only the original $p$ points are tiling.
Now since $A$ is completely inside the subgroup spanned by $e_{1},\ldots,e_{p-1}$, it suffices for us to look at the case at d=p1d=p-1 only. Then by Lemma 7 we easily see that $A$ is indeed a tiling set in $\mathbb{Z}_{p}^{p-1}$, thus also in $\mathbb{Z}^{p-1}$. ∎
The condition that $p$ points have to be in general linear positions is tight in the sense that $3$ colinear points need not be tiling, for example, $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1). The author does not know non-tiling examples for other values of $p$, but would expect that they exist for every pair $p,d$ with d<p1d<p-1.
Acknowldgement
The author would like to thank Tao Zhang (XDU) for the hospitality in Xi’an, and the Chern Institute of Mathematics (Tianjin) for hosting while preparing for this paper.
The author thanks Jia Zheng (WHU) for mentioning the phenomenon described in Theorem 2 at $p=3$ to him, as well as Shilei Fan (CCNU) and Gergely Kiss (ARIM) for short communications that improved Lemma 5.

### v3

@@
 Fuglede conjectured that being spectral and being tiling are actually equivalent [6], this has been studied extensively and disproved in general when the dimension is at least three [5, 10, 11, 17, 21] but remains open for one dimensional and two dimensional cases.
-Researches over the Fuglede conjecture in finite Abelian groups as well as in $p$-adic fields have been active in the past years, see e.g., [3, 4, 9, 12, 13, 15, 16, 21] (the full list is of course too long to be enumerated here, this is only an attemp to mention some of the results in one dimensional or two dimensional cases, see perhaps references therein for more on the history and development of this problem).
+Researches over the Fuglede conjecture in finite Abelian groups as well as in $p$-adic fields have been active in the past years, see e.g., [3, 4, 9, 12, 13, 15, 16, 22] (the full list is of course too long to be enumerated here, this is only an attempt to mention some of the results in one dimensional or two dimensional cases, see perhaps references therein for more on the history and development of this problem).
 It is well known that a tiling set of prime size in $\mathbb{Z}^{d}$ must have at least one periodic tiling complement [20] (it may be worthing mentioning that being periodic in $d$ dimensions means having $d$ linearly independent periods, while having periods in one direction is called weakly periodic instead, readers may also find the periodic tiling conjecture and its positive and negative cases in e.g., [1, 7, 8, 14] relevant).
@@
 Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$.
-It is also worth mentioning that for $d\geq p$ the second theorem is acutually already included in [17, Corollary 2.4].
+It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [17, Corollary 2.4].
 The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$.
@@
 $\widehat{\mathbf{1}_{A}}(\xi)=\sum_{a\in A}\operatorname{mul}(a)\cdot e^{2\pi i\langle a,\xi\rangle/n},$
-where $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. It is straigtforward to verify that
+where $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. It is straightforward to verify that
 […]
@@
 Proof.
-Assume the contrary that $\mathbf{1}_{A}$ is not a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, then let $m$ be the smallest multipicity of elements in $A$, i.e.,
+Assume the contrary that $\mathbf{1}_{A}$ is not a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, then let $m$ be the smallest multiplicity of elements in $A$, i.e.,
 $m=\min\{\operatorname{mul}(a):a\in A\},$
@@
 Let $p$ be a prime number and $A\subset\mathbb{Z}^{d}$ with $|A|=p$, if $A\oplus B=\mathbb{Z}^{d}$, then there is some $n\in\mathbb{N}$ and $B^{\prime}\in\mathbb{Z}_{n}^{d}$ such that $A\mapsto\pi_{n}(A)$ is injective and $\pi_{n}(A)\oplus B^{\prime}=\mathbb{Z}_{n}^{d}$.
-The way that Lemma 2 is stated here is slightly different from its original version in [19, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.
+The way that Lemma 2 is stated here is slightly different from its original version in [20, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.
 Lemma 3.
@@
 Lemma 4.
-Let $E$ be an equaivlence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $A$ annihilates one element of $E$, then it annihilates entire $E$.
+Let $E$ be an equivalence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $A$ annihilates one element of $E$, then it annihilates entire $E$.
 Proof.
@@
 Lemma 5.
-Let $E$ be an equaivlence class under $\sim$ in some finite Abelian group. If $p$ is the smallest prime divisor of $\operatorname{ord}(E)$, then $E$ admits a derived set of size $p$.
+Let $E$ be an equivalence class under $\sim$ in some finite Abelian group. If $p$ is the smallest prime divisor of $\operatorname{ord}(E)$, then $E$ admits a derived set of size $p$.
 Proof.
@@
 Repeating the argument on every member of $\Delta\mathbb{Z}_{p^{k}}^{d}$ leads to (6).
-Now as $\pi_{p^{k}}(B^{\prime})$ is a multiset on $\mathbb{Z}_{p^{k}}^{d}$, Lemma 1 indicates that $\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}$ is just a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{p^{k}}^{d}}$, which means $|Z_{p^{k}}^{d}|$ divides $|\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}|$.
-Since $|\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}|=|B^{\prime}|$, this further indicates that $p^{kd}$ divides $|B^{\prime}|$, consequently we would have $p^{kd}m^{d}=|\mathbb{Z}_{n}^{d}|=|A|\cdot|B|$ is divisible by $p^{kd+1}$, which is a contradiction.
+Now as $\pi_{p^{k}}(B^{\prime})$ is a multiset on $\mathbb{Z}_{p^{k}}^{d}$, Lemma 1 indicates that $\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}$ is just a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{p^{k}}^{d}}$, which means $|Z_{p^{k}}^{d}|$ divides $|\pi_{p^{k}}(B^{\prime})|$.
+Since $|\pi_{p^{k}}(B^{\prime})|=|B^{\prime}|$, this further indicates that $p^{kd}$ divides $|B^{\prime}|$, consequently we would have $p^{kd}m^{d}=|\mathbb{Z}_{n}^{d}|=|A|\cdot|B|$ is divisible by $p^{kd+1}$, which is a contradiction.
 ∎
 Lemma 6.
-Given any $n$ linearly indpenedent vectors $v_{1},\ldots,v_{n}$ on $\mathbb{Z}^{n}$, there exists a group automorphism of the additive group $\mathbb{Z}^{n}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,1,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,1)$.
+Given any $n$ linearly indenpendent vectors $v_{1},\ldots,v_{n}$ on $\mathbb{Z}^{n}$, there exists a group automorphism of the additive group $\mathbb{Z}^{n}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,1,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,1)$.
 In particular, if $p$ is a prime number and n=p1n=p-1, then there exists a group automorphism of the additive group $\mathbb{Z}^{p-1}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$.
@@
 $\mathbb{Z}^{n}\xrightarrow{\qquad\qquad V^{\prime}\qquad\qquad}D\mathbb{Z}^{n}\xrightarrow{\qquad x\mapsto x/D\qquad}\mathbb{Z}^{n},$
-If n=p1n=p-1, then $\operatorname{diag}(1,2,\ldots,{p-1})$ is also a group autophism on $\mathbb{Z}^{p-1}$, thus the desired automorphism is constructed by a further composition:
-$\mathbb{Z}^{p-1}\xrightarrow{\qquad\qquad\qquad V^{\prime}\qquad\qquad\qquad}D\mathbb{Z}^{p-1}\xrightarrow{\qquad\qquad x\mapsto x/D\qquad\qquad}\mathbb{Z}^{p-1}\xrightarrow{\qquad\operatorname{diag}(1,2,\ldots,{p-1})\qquad}\mathbb{Z}^{p-1}.$
+If n=p1n=p-1, then $\operatorname{diag}(1,2,\ldots,{p-1})$ is also a group automorphism on $\mathbb{Z}_{p}^{p-1}$, thus the desired automorphism is constructed by a further composition:
+$\mathbb{Z}^{p-1}\xrightarrow{\qquad V^{\prime}\qquad}D\mathbb{Z}^{p-1}\xrightarrow{\;x\mapsto x/D\;}\mathbb{Z}^{p-1}\xrightarrow{\;\operatorname{diag}(1,2,\ldots,{p-1})\otimes I\;}\mathbb{Z}^{p-1}\left(=\mathbb{Z}_{p}^{p-1}\oplus(p\mathbb{Z})^{p-1}\right),$
+where $I$ is identity so that $\operatorname{diag}(1,2,\ldots,{p-1})$ only acts on the $\mathbb{Z}_{p}^{p-1}$ part. ∎
 Lemma 7.
@@
 then it will follow from (2) that $(A,B)$ forms a tiling pair.
-Since $p$ is a prime number, the (p1)(p-1)-st roots of unity in the finite field $\mathbb{F}_{p}$ is precisely $1,\ldots,p-1$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and its columns, denoted by $u_{1},\ldots,u_{p-1}$ is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$).
+Since $p$ is a prime number, the (p1)(p-1)-st roots of unity in the finite field $\mathbb{F}_{p}$ is precisely $1,\ldots,p-1$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and the set of its columns, denoted by $u_{1},\ldots,u_{p-1}$, is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$).
 In particular, since $x$ is a column of it we may without loss of generality take $u_{1}=x$.
@@
 ∎
-The condition that $p$ points have to be in general linear positions is tight in the sense that $3$ colinear points need not be tiling, for example, $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1).
+The condition that $p$ points have to be in general linear positions is tight in the sense that $3$ colinear points need not be tiling, for example, $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [18] for a characterization of prime power size tiles in $\mathbb{Z}$).
 The author does not know non-tiling examples for other values of $p$, but would expect that they exist for every pair $p,d$ with d<p1d<p-1.

### v4

@@
 Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$.
-It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [17, Corollary 2.4].
+It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [18, Corollary 2.4].
 The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$.
 The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, while the existence of such a set being included in the zero set of $A$ is shown by contradiction.
-Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$, thus leading to a contradiction if $A$ does not admit such a spectrum.
+Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$ (a statement that is similar to the Coven-Meyerowitz T1 property in [2]), thus leading to a contradiction if $A$ does not admit such a spectrum.
 The proof of the second theorem then reduces to proving any $p$ points in general linear positions are tiling, for which we first apply a group isomorphism to map them to the origin and $(1,0\ldots,0),(0,2,0,\ldots,0),\ldots,(0,\ldots,0,p-1)$, then we construct a tiling complement for such a set explicitly.
@@
 Let $p$ be a prime number and $A\subset\mathbb{Z}^{d}$ with $|A|=p$, if $A\oplus B=\mathbb{Z}^{d}$, then there is some $n\in\mathbb{N}$ and $B^{\prime}\in\mathbb{Z}_{n}^{d}$ such that $A\mapsto\pi_{n}(A)$ is injective and $\pi_{n}(A)\oplus B^{\prime}=\mathbb{Z}_{n}^{d}$.
-The way that Lemma 2 is stated here is slightly different from its original version in [20, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.
+The way that Lemma 2 is stated here is slightly different from its original version in [21, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.
 Lemma 3.

### v5

@@
 Any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
-Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$.
+Here $p$ points being in general linear positions means that they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$.
 It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [18, Corollary 2.4].
@@
 Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$ (a statement that is similar to the Coven-Meyerowitz T1 property in [2]), thus leading to a contradiction if $A$ does not admit such a spectrum.
-The proof of the second theorem then reduces to proving any $p$ points in general linear positions are tiling, for which we first apply a group isomorphism to map them to the origin and $(1,0\ldots,0),(0,2,0,\ldots,0),\ldots,(0,\ldots,0,p-1)$, then we construct a tiling complement for such a set explicitly.
+The proof of the second theorem then reduces to proving any $p$ points in general linear positions are tiling, for which we can simply show that they tile the sublattice they generate.
+This follows from the fact that the set formed by the origin and $(1,0\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$ is tiling (a tiling complement can be found explicitly by computation), then we can map this set to the $p$ points we are looking at by a linear transform.
 2 Preliminaries
@@
 is its orthogonal group in $\mathbb{Z}_{n}^{d}$.
-That(3) holds for cyclic subgroups is quite trivial, while non-cyclic subgroups can be viewed as products of cyclic ones, its orthogonal set is then the intersection over orthogonal sets of its cyclic factors.
+That (3) holds for cyclic subgroups is quite trivial, while non-cyclic subgroups can be viewed as products of cyclic ones, then its orthogonal set is simply the intersection over orthogonal sets of its cyclic factors.
 Let the support of a function $f$ in $\mathbb{Z}_{n}^{d}$ be defined as
@@
 Lemma 6.
-Given any $n$ linearly indenpendent vectors $v_{1},\ldots,v_{n}$ on $\mathbb{Z}^{n}$, there exists a group automorphism of the additive group $\mathbb{Z}^{n}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,1,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,1)$.
-In particular, if $p$ is a prime number and n=p1n=p-1, then there exists a group automorphism of the additive group $\mathbb{Z}^{p-1}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$.
-Proof.
-Let $V$ be the matrix whose columns are $v_{1},\ldots,v_{n}$, and $D$ the determinant of $V$. Clearly $D\neq 0$ because of linear independence.
-Let $V^{\prime}$ be the adjugate matrix of $V$, i.e., if $U_{i,j}$ is the determinant of the matrix obtained by deleting the $i$-th row and the $j$-th column of $V$, then $V^{\prime}$ is the transpose of the matrix whose $i,j$-th entry is $(-1)^{i+j}U_{i,j}$.
-Then $V^{\prime}V=DI$, and thus $V^{\prime}$ is a group isomorphism from $\mathbb{Z}^{n}$ to $D\mathbb{Z}^{n}$ that maps $v_{1},\ldots,v_{p-1}$ to $(D,0,\ldots,0)$, $(0,D,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,D)$.
-The desired automorphism can then be constructed by composing the following two isomorphisms:
-$\mathbb{Z}^{n}\xrightarrow{\qquad\qquad V^{\prime}\qquad\qquad}D\mathbb{Z}^{n}\xrightarrow{\qquad x\mapsto x/D\qquad}\mathbb{Z}^{n},$
-If n=p1n=p-1, then $\operatorname{diag}(1,2,\ldots,{p-1})$ is also a group automorphism on $\mathbb{Z}_{p}^{p-1}$, thus the desired automorphism is constructed by a further composition:
-$\mathbb{Z}^{p-1}\xrightarrow{\qquad V^{\prime}\qquad}D\mathbb{Z}^{p-1}\xrightarrow{\;x\mapsto x/D\;}\mathbb{Z}^{p-1}\xrightarrow{\;\operatorname{diag}(1,2,\ldots,{p-1})\otimes I\;}\mathbb{Z}^{p-1}\left(=\mathbb{Z}_{p}^{p-1}\oplus(p\mathbb{Z})^{p-1}\right),$
-where $I$ is identity so that $\operatorname{diag}(1,2,\ldots,{p-1})$ only acts on the $\mathbb{Z}_{p}^{p-1}$ part. ∎
-Lemma 7.
-The set
-$A=\{(0,0,\ldots,0),(1,0,\ldots,0),(0,2,0,\ldots,0),\ldots,(0,\ldots,0,p-1)\},$
+Let $e_{i}$ be the $i$-th standard Euclidean basis, then the set
+$A=\{0,e_{1},2e_{2},\ldots,(p-1)e_{p-1}\},$
 is both tiling and spectral in $\mathbb{Z}_{p}^{p-1}$.
 Proof.
-Let $x=(1,\ldots,1)$, and $\omega=e^{2\pi i/p}$, then it is easy to verify that
+Let $x=(1,\ldots,1)$, and set $\omega=e^{2\pi i/p}$, then it is easy to verify that
 $\widehat{\mathbf{1}_{A}}(x)=1+\omega+\ldots+\omega^{p-1}=0.$
-Hence the cyclic group generated by $x$ (denoted by $\langle x\rangle$) is a spectrum of $A$. To see that $A$ is tiling, we will construct a set $B$ so that $|A|\cdot|B|=|\mathbb{Z}_{p}^{p-1}|$ and
+Therefore the cyclic group generated by $x$ (denoted by $\langle x\rangle$) is a spectrum of $A$.
+To see that $A$ is tiling, we will construct a set $B$ so that $|A|\cdot|B|=|\mathbb{Z}_{p}^{p-1}|$ and
 $Z(\widehat{\mathbf{1}_{B}})=\mathbb{Z}_{p}^{p-1}\setminus\langle x\rangle,$ (8)
 then it will follow from (2) that $(A,B)$ forms a tiling pair.
-Since $p$ is a prime number, the (p1)(p-1)-st roots of unity in the finite field $\mathbb{F}_{p}$ is precisely $1,\ldots,p-1$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and the set of its columns, denoted by $u_{1},\ldots,u_{p-1}$, is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$).
-In particular, since $x$ is a column of it we may without loss of generality take $u_{1}=x$.
+Since $p$ is a prime number, each of $1,\ldots,p-1$ is $a$ (p1)(p-1)-st root of unity in the finite field $\mathbb{F}_{p}$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and the set of its columns, denoted by $u_{1},\ldots,u_{p-1}$, is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$).
+In particular, since $x$ is a column of it we may without loss of generality order the columns in the way so that $u_{1}=x$.
 Let $B$ be the span (over $\mathbb{F}_{p}$) of $u_{2},\ldots,u_{p-1}$, then $|B|=p^{p-2}$ is of the correct size, and we simply notice that $B$ and $\langle x\rangle$ are orthogonal groups of each other, thus (8) follows immediately from the Poisson summation formula (3).
+∎
+Corollary 1.
+The set
+$A=\{0,e_{1},\ldots,e_{p-1}\},$
+is both tiling and spectral in $\mathbb{Z}^{p-1}$.
+Proof.
+It suffices to show that $A$ is tiling in $\mathbb{Z}_{p}^{p-1}$ since if $(A,B)$ is a tiling pair in $\mathbb{Z}_{p}^{p-1}$, then $(A,B\oplus(p\mathbb{Z})^{p-1})$ becomes a tiling pair in $\mathbb{Z}^{p-1}$.
+Let
+$S=\{0,e_{1},2e_{2},\ldots,(p-1)e_{p-1}\},$
+be the set defined in Lemma 6, and set $D=\operatorname{diag}(1,2,\ldots,{p-1})$.
+Then $D$ is a group automorphism on $\mathbb{Z}_{p}^{p-1}$, its inverse is simply $D^{-1}=\operatorname{diag}(1,2^{-1},\ldots,(p-1)^{-1})$ where for each $k\in\{1,\ldots,p-1\}$, $k^{-1}$ is the multiplicative inverse of $k$ in $\mathbb{Z}_{p}$, each $k^{-1}$ exists since $p$ is a prime number.
+By Lemma 6, $S$ is tiling in $\mathbb{Z}_{p}^{p-1}$, therefore $A=D^{-1}S$ is also tiling in $D^{-1}\mathbb{Z}_{p}^{p-1}$, but $D^{-1}\mathbb{Z}_{p}^{p-1}$ equals $\mathbb{Z}_{p}^{p-1}$, thus $A$ is indeed tiling in $\mathbb{Z}_{p}^{p-1}$.
+Its spectrality then follows from Theorem 1.
 ∎
@@
 In light of Theorem 1, it suffices to show that these points are tiling in $\mathbb{Z}^{d}$.
-Without loss of generality we may assume that one of the point is the origin, and the other p1p-1 points are $v_{1},\ldots,v_{p-1}$.
-By Lemma 6, we can find a group automorphism of $\mathbb{Z}^{d}$ that maps the original $p$ points $0,v_{1},\ldots,v_{p-1}$ to
-$A=\{0,e_{1},2e_{2},\ldots,(p-1)e_{p-1}\},$
-where $e_{i}$ is the $i$-th standard Euclidean basis. Since it is a group automorphism we assert that $A$ is tiling if and only the original $p$ points are tiling.
-Now since $A$ is completely inside the subgroup spanned by $e_{1},\ldots,e_{p-1}$, it suffices for us to look at the case at d=p1d=p-1 only.
-Then by Lemma 7 we easily see that $A$ is indeed a tiling set in $\mathbb{Z}_{p}^{p-1}$, thus also in $\mathbb{Z}^{p-1}$.
+Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$. Let $V$ be the linear transform defined as
+$Ve_{i}=\begin{cases}v_{i},&\text{if }i=1,\ldots,p-1;\\ e_{i},&\text{if }i>p-1.\end{cases}$
+$V$ has full rank since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions. Further set
+$A=\{0,v_{1},\ldots,v_{p-1}\},\qquad E=\{0,e_{1},\ldots,e_{p-1}\}.$
+By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the space obtained by applying $V$ to $\operatorname{span}(E)$, which is a sublattice since $\operatorname{span}(E)=\mathbb{Z}^{p-1}$.
+This implies that $A$ is tiling in $\mathbb{Z}^{d}$, and its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$.
 ∎
-The condition that $p$ points have to be in general linear positions is tight in the sense that $3$ colinear points need not be tiling, for example, $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [19] for a characterization of prime power size tiles in $\mathbb{Z}$).
-The author does not know non-tiling examples for other values of $p$, but would expect that they exist for every pair $p,d$ with d<p1d<p-1.
+Theorem 2 produces some interesting consequences, for example we can assert that any $3$ points must be tiling as long as they are not on a line.
+On the other hand, this also shows that the condition that $p$ points have to be in general linear positions is tight, since $3$ colinear points indeed need not be tiling, e.g., $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [19] for a characterization of prime power size tiles in $\mathbb{Z}$).
 Acknowldgement

### v6

@@
 In light of Theorem 1, it suffices to show that these points are tiling in $\mathbb{Z}^{d}$.
-Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$. Let $V$ be the linear transform defined as
-$Ve_{i}=\begin{cases}v_{i},&\text{if }i=1,\ldots,p-1;\\ e_{i},&\text{if }i>p-1.\end{cases}$
-$V$ has full rank since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions. Further set
+Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$.
+Let $V$ be an arbitrary linear transform that maps $e_{i}$ to $v_{i}$ for i<=p1i<=p-1, such $V$ exists since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions (which means that $v_{1},\ldots,v_{p-1}$ are linearly independent).
+Further set
 $A=\{0,v_{1},\ldots,v_{p-1}\},\qquad E=\{0,e_{1},\ldots,e_{p-1}\}.$
 By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the space obtained by applying $V$ to $\operatorname{span}(E)$, which is a sublattice since $\operatorname{span}(E)=\mathbb{Z}^{p-1}$.
-This implies that $A$ is tiling in $\mathbb{Z}^{d}$, and its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$.
+This implies that $A$ is also tiling in $\mathbb{Z}^{d}$.
+Its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$.
 ∎

### v7

@@
 The proof is by contradiction, it is simply shown that the tiling complement of such a tile can not annihilate all $p$-subgroups.
-In addition, with a simple transformation we prove that any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
+In addition, with a simple transformation we prove that any $p$ points in general linear positions in $\mathbb{Z}^{d}(d\geq p-1)$ must be both tiling and spectral.
 Keywords: Fuglede conjecture; tiling sets; spectral sets.
@@
 1 Introduction
-Let $G$ be a finitely generated Abelian group, a set $A$ is tiling (or is a tile) in $G$ if there exists some $B$ such that $\{A+b\}_{b\in B}$ forms a partition of $G$, and is spectral if there exists a set of group characters of $G$ that forms an orthogonal basis on $L^{2}(A)$.
+Let $G$ be a finitely generated Abelian group, a set $A$ is tiling (or is a tile) in $G$ if there exists some $B$ such that $\{A+b\}_{b\in B}$ forms a partition of $G$, and is spectral if there exists a set $S$ of group characters of $G$ that forms an orthogonal basis on $L^{2}(A)$.
 In such cases $B$ (resp.
@@
 Theorem.
-Any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
-Here $p$ points being in general linear positions means that they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$.
-It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [18, Corollary 2.4].
+If $p$ is a prime number, then any $p$ points in general linear positions in $\mathbb{Z}^{d}(d\geq p-1)$ must be both tiling and spectral.
+Here $p$ points being in general linear positions means that they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$, therefore this necessarily requires $d\geq p-1$.
+The second theorem leads to some interesting findings, for example, it follows that any $3$ points in $\mathbb{Z}^{d}(d>1)$ must be both spectral and tiling if they are not on the same line.
+It is also worth mentioning that $p$ linearly independent points must be tiling is already included in [18, Corollary 2.4].
 The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$.
-The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, while the existence of such a set being included in the zero set of $A$ is shown by contradiction.
-Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$ (a statement that is similar to the Coven-Meyerowitz T1 property in [2]), thus leading to a contradiction if $A$ does not admit such a spectrum.
+The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, from each equivalence class we can squeeze out a derived set of size $p$ whose difference set is still in the same class.
+The existence of a spectrum within one of these classes is then shown by contradiction: Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$ (a statement that is similar to the Coven-Meyerowitz T1 property in [2]), thus leading to a contradiction if $A$ does not admit a spectrum.
 The proof of the second theorem then reduces to proving any $p$ points in general linear positions are tiling, for which we can simply show that they tile the sublattice they generate.
@@
 2 Preliminaries
-Let $A$ be a multiset (i.e., a set that allows repetitive elements) on $\mathbb{Z}_{n}^{d}$ (i.e., each element of $A$ is a member of $\mathbb{Z}_{n}^{d}$), the multiplicity of an element $a\in A$ (i.e., the number of copies of $a$ in $A$) will be denoted by $\operatorname{mul}(a)$, the characteristic function on such a multiset $A$ will be defined and written as
-$\mathbf{1}_{A}(x)=\begin{cases}\operatorname{mul}(x)&x\in A,\\ 0&x\notin A.\end{cases}$
-The Fourier transform of $\mathbf{1}_{A}(x)$ on $\mathbb{Z}_{n}^{d}$ is
-$\widehat{\mathbf{1}_{A}}(\xi)=\sum_{a\in A}\operatorname{mul}(a)\cdot e^{2\pi i\langle a,\xi\rangle/n},$
-where $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. It is straightforward to verify that
-[…]
-[…]
-where the difference set $\Delta S$ of a usual set (i.e., a set with distinct elements only) $S$ and the zero set $Z(f)$ of a function $f$ is defined respectively as
-$\Delta S=\{s-s^{\prime}:s,s^{\prime}\in S,\;s\neq s^{\prime}\},\qquad Z(f)=\{x:f(x)=0\}.$
+Let $f$ be a function on $\mathbb{Z}_{n}^{d}$, then its Fourier transform on $\mathbb{Z}_{n}^{d}$ is
+$\hat{f}(\xi)=\sum_{x\in\mathbb{Z}_{n}^{d}}f(x)\cdot e^{2\pi i\langle x,\xi\rangle/n},\qquad\xi\in\mathbb{Z}_{n}^{d}.$
+Denote by
+$\operatorname{supp}(f)=\{x\in\mathbb{Z}_{n}^{d}:f(x)\neq 0\},\qquad Z(f)=\{x\in\mathbb{Z}_{n}^{d}:f(x)=0\},$
+the support and the zero set of $f$ respectively, then the uncertainty principle on finite Abelian groups [3, 20] asserts that
+$|\operatorname{supp}(f)|\cdot|\operatorname{supp}(\widehat{f})|\geq|\mathbb{Z}_{n}^{d}|,$ (1)
+holds for any non-zero function $f$ defined on $\mathbb{Z}_{n}^{d}$.
+If $A$ is a finite multiset (i.e., a set that allows repetitive elements) on $\mathbb{Z}_{n}^{d}$ (i.e., each element of $A$ is a member of $\mathbb{Z}_{n}^{d}$), and the multiplicity of an element $a\in A$ (i.e., the number of copies of $a$ in $A$) is denoted by $m_{a}$, then the characteristic function on such a multiset $A$ can be defined and written as
+$\mathbf{1}_{A}(x)=\begin{cases}m_{x}&x\in A,\\ 0&x\notin A.\end{cases}$
+Its Fourier transform on $\mathbb{Z}_{n}^{d}$ is
+$\widehat{\mathbf{1}}_{A}(\xi)=\sum_{a\in A}m_{a}\cdot e^{2\pi i\langle a,\xi\rangle/n}.$
+If $A,B$ are finite multisets, and $A+B$ is the multiset obtained by enumerating $a\in A,b\in B$ (so $a$ appears $m_{a}$ times and $b$ appears $m_{b}$ times) and take the collection of all sums $a+b$, then straightforward computation shows
+$\widehat{\mathbf{1}}_{A+B}=\widehat{\mathbf{1}}_{A}\cdot\widehat{\mathbf{1}}_{B}.$
+With these settings we easily get that when $A,B,S$ are usual sets (every element appear only once), then
+$(A,S)\text{ is a spectral pair on }\mathbb{Z}_{n}^{d}$ $\qquad\Leftrightarrow\qquad\begin{cases}|A|=|S|,\\ \Delta S\subseteq Z(\widehat{\mathbf{1}}_{A}),\end{cases}$ (2)
+$(A,B)\text{ is a tiling pair on }\mathbb{Z}_{n}^{d}$ $\qquad\Leftrightarrow\qquad\begin{cases}|A|\cdot|B|=n^{d},\\ \Delta\mathbb{Z}_{n}^{d}\subseteq Z(\widehat{\mathbf{1}}_{A})\cup Z(\widehat{\mathbf{1}}_{B}),\end{cases}$ (3)
+where the difference set $\Delta S$ is defined as
+$\Delta S=\{s-s^{\prime}:s,s^{\prime}\in S,\;s\neq s^{\prime}\}.$
 One may also verify the Poisson summation formula below:
@@
 is its orthogonal group in $\mathbb{Z}_{n}^{d}$.
-That (3) holds for cyclic subgroups is quite trivial, while non-cyclic subgroups can be viewed as products of cyclic ones, then its orthogonal set is simply the intersection over orthogonal sets of its cyclic factors.
-Let the support of a function $f$ in $\mathbb{Z}_{n}^{d}$ be defined as
-$\operatorname{supp}(f)=\mathbb{Z}_{n}^{d}\setminus Z(f),$
-then the uncertainty principle on finite Abelian groups [3, 20] asserts that
-$|\operatorname{supp}(f)|\cdot|\operatorname{supp}(\widehat{f})|\geq|\mathbb{Z}_{n}^{d}|,$ (4)
-holds for any non-zero function $f$ defined on $\mathbb{Z}_{n}^{d}$.
+(4) can be first established on cyclic subgroups (which is quite trivial), then since non-cyclic subgroups are products of cyclic ones, its orthogonal set is simply the intersection over orthogonal sets of its cyclic factors.
 Lemma 1.
-Let $A$ be a multiset on $\mathbb{Z}_{n}^{d}$, if $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{A})$, then $\mathbf{1}_{A}$ must be a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, i.e., $\mathbf{1}_{A}=m\mathbf{1}_{\mathbb{Z}_{n}^{d}}$ for some $m\in\mathbb{N}$.
+Let $A$ be a multiset on $\mathbb{Z}_{n}^{d}$, if $\mathbb{Z}_{n}^{d}\setminus\{0\}\subseteq Z(\widehat{\mathbf{1}}_{A})$, then $\mathbf{1}_{A}$ must be a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, i.e., $\mathbf{1}_{A}=m\mathbf{1}_{\mathbb{Z}_{n}^{d}}$ for some $m\in\mathbb{N}$.
 Proof.
 Assume the contrary that $\mathbf{1}_{A}$ is not a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, then let $m$ be the smallest multiplicity of elements in $A$, i.e.,
-$m=\min\{\operatorname{mul}(a):a\in A\},$
+$m=\min\{m_{a}:a\in A\},$
 and consider
@@
 On the other hand, it is straightforward to check that
-$Z(\widehat{\mathbf{1}_{\mathbb{Z}_{n}^{d}}})=\mathbb{Z}_{n}^{d}\setminus\{0\}=\Delta\mathbb{Z}_{n}^{d},$
-i.e., $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{\mathbb{Z}_{n}^{d}})$, thus combined with the assumption that $\Delta\mathbb{Z}_{n}^{d}\subseteq Z(\mathbf{1}_{A})$ we get
+$Z(\widehat{\mathbf{1}}_{\mathbb{Z}_{n}^{d}})=\mathbb{Z}_{n}^{d}\setminus\{0\},$
+thus combined with the assumption that $\mathbb{Z}_{n}^{d}\setminus\{0\}\subseteq Z(\widehat{\mathbf{1}}_{A})$ we get
 […]
@@
 so that $\pi_{n}$ is the projection from $\mathbb{Z}^{d}$ to $\mathbb{Z}_{n}^{d}$ (as a map between sets).
-Given a set $A\subseteq\mathbb{Z}^{d}$, $\pi_{n}(A)$ will denote the multiset obtained by applying $\pi_{n}$ on every element of $A$.
+Given a finite set $A\subseteq\mathbb{Z}^{d}$, $\pi_{n}(A)$ will denote the multiset obtained by applying $\pi_{n}$ on every element of $A$.
 Clearly $\pi_{n}(A)$ will only be a usual set if (and only if) the map $A\mapsto\pi_{n}(A)$ is injective.
@@
 Lemma 3.
-Let $n=p^{k}m$ with $p$ being a prime number that does not divide $m$, and set $A^{\prime}=\pi_{p^{k}}(A)$.
-Let both $f,g$ be characteristic functions on $A^{\prime}$ but with the domain of $f$ being the entire $\mathbb{Z}_{n}^{d}$ while the domain of $g$ being $\mathbb{Z}_{p^{k}}^{d}$ only.
-If $x\in\mathbb{Z}_{n}^{d}$ is an element of order $p^{t}$ for some $1\leq t\leq k$ , then we have
+Let $n=p^{k}m$ with $p$ being a prime number that does not divide $m$, and $A\subseteq\mathbb{Z}_{n}^{d}$.
+Set $A^{\prime}=\pi_{p^{k}}(A)$, and let $f,g$ both be characteristic functions on $A^{\prime}$, but defined on $\mathbb{Z}_{n}^{d}$ and $\mathbb{Z}_{p^{k}}^{d}$ respectively so that the Fourier transforms
+$\widehat{\mathbf{1}}_{A}(x)=\sum_{a\in A}e^{2\pi i\langle a,\xi\rangle/n},\qquad\hat{f}(\xi)=\sum_{a^{\prime}\in A^{\prime}}m_{a^{\prime}}e^{2\pi i\langle a^{\prime},\xi\rangle/n},\qquad\xi\in\mathbb{Z}_{n}^{d},$
+take place in $\mathbb{Z}_{n}^{d}$, while
+$\hat{g}(\xi^{\prime})=\sum_{a^{\prime}\in A^{\prime}}m_{a^{\prime}}e^{2\pi i\langle a^{\prime},\xi^{\prime}\rangle/p^{k}},\qquad\xi^{\prime}\in\mathbb{Z}_{p^{k}}^{d},$
+is performed in $\mathbb{Z}_{p^{k}}^{d}$.
+If $x=(x_{1},\ldots,x_{d})\in\mathbb{Z}_{n}^{d}$ is an element of order $p^{t}$ for some $1\leq t\leq k$ (which means every coordinate $x_{1},\ldots,x_{d}$ must be divisible by $m$), then we have
 $\widehat{\mathbf{1}}_{A}(x)=\widehat{f}(x)=\widehat{g}(x^{\prime}),$
-where $x^{\prime}\in\mathbb{Z}_{p^{k}}^{d}$ is the element that satisfies $x=mx^{\prime}$, and the last Fourier transform $\hat{g}$ is performed in $\mathbb{Z}_{p^{k}}^{d}$ only (the first two Fourier transforms $\widehat{\mathbf{1}_{A}}$ and $\widehat{f}$ take place in $\mathbb{Z}_{n}^{d}$).
+where $x^{\prime}=(x_{1}/m,\ldots,x_{d}/m)\in\mathbb{Z}_{p^{k}}^{d}$.
 Proof.
@@
 $A^{\prime}=\pi_{p^{k}}(A)=\{r_{a}:a\in A\}.$
-The support set of $f,g$ coincide, but they are defined on different subgroups, thus we have
-$\widehat{f}(\xi)=\sum_{a\in A}e^{2\pi i\langle r_{a},\xi\rangle/n},\qquad\widehat{g}(\xi)=\sum_{a\in A}e^{2\pi i\langle r_{a},\xi\rangle/p^{k}}$
-in which Fourier transforms in $\mathbb{Z}_{n}^{d}$ and in $\mathbb{Z}_{p^{k}}^{d}$ are applied respectively.
 Then direct computation shows that
@@
 Lemma 4.
-Let $E$ be an equivalence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $A$ annihilates one element of $E$, then it annihilates entire $E$.
+Let $E$ be an equivalence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$.
+If $Z(\widehat{\mathbf{1}}_{A})\cap E\neq\emptyset$, then $E$ is completely contained in $Z(\widehat{\mathbf{1}}_{A})$, i.e., $E\subseteq Z(\widehat{\mathbf{1}}_{A})$.
 Proof.
 Set $m=\operatorname{ord}(E)$ and $r=n/m$. Let $x$ be an arbitrary element of $E$, then $x=rx^{\prime}$ for some $x^{\prime}\in\mathbb{Z}_{m}^{d}$, consequently
-$\widehat{1_{A}}(x)=\sum_{a\in A}e^{2\pi i\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\langle a,x^{\prime}\rangle/m}.$
+$\widehat{\mathbf{1}}_{A}(x)=\sum_{a\in A}e^{2\pi i\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\langle a,x^{\prime}\rangle/m}.$
 We shall view the above expression as the polynomial
@@
 evaluated at the $m$-th root of unity $\omega=e^{2\pi i/m}$, i.e.,
-$\widehat{1_{A}}(x)=P(\omega).$
+$\widehat{\mathbf{1}}_{A}(x)=P(\omega).$
 […]
-$\widehat{1_{A}}(y)=\sum_{a\in A}e^{2\pi i\langle a,\lambda x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x^{\prime}\rangle/m}=P(\omega^{\lambda})=0,$
+$\widehat{\mathbf{1}}_{A}(y)=\sum_{a\in A}e^{2\pi i\langle a,\lambda x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x\rangle/n}=\sum_{a\in A}e^{2\pi i\lambda\langle a,x^{\prime}\rangle/m}=P(\omega^{\lambda})=0,$
 where the last equality holds since $\lambda$ is coprime to $m$, and so that $\omega^{\lambda}$ is also a root of $\Phi_{m}(z)$. This finishes the proof since $x,y$ are chosen arbitrarily from $E$. ∎
@@
 Let $m$ be the number so that $n=p^{k}m$ and $\gcd(p,m)=1$.
-[…]
+Consider all equivalence classes in $\mathbb{Z}_{n}^{d}$ whose orders are powers of $p$, by Lemma 4, each equivalence class must be either completely inside $Z(\widehat{\mathbf{1}}_{A^{\prime}})$ or completely disjoint with $Z(\widehat{\mathbf{1}}_{A^{\prime}})$.
 By Lemma 5, each of these equivalence classes has a derived set of size $p$, thus if any of these classes is in $Z(\widehat{\mathbf{1}}_{A^{\prime}})$, then the corresponding derived set would be a spectrum of $A^{\prime}$ and we are done.
@@
 By Lemma 3 this further implies
-[…]
+$\Delta\mathbb{Z}_{p^{k}}^{d}\subseteq Z(\widehat{\mathbf{1}}_{\pi_{p^{k}}(B^{\prime})}).$ (6)
 Indeed, every element $x^{\prime}\in\Delta\mathbb{Z}_{p^{k}}^{d}$ can be lifted to an element of the same order in $\mathbb{Z}_{n}^{d}$ through the map $x^{\prime}\mapsto mx^{\prime}$, and Lemma 3 asserts that
@@
 Theorem 2.
-Any $p$ points in general linear positions in $\mathbb{Z}^{d}$ must be both tiling and spectral if $d\geq p-1$.
+If $p$ is a prime number, then any $p$ points in general linear positions in $\mathbb{Z}^{d}(d\geq p-1)$ must be both tiling and spectral.
 Proof.
@@
 Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$.
-Let $V$ be an arbitrary linear transform that maps $e_{i}$ to $v_{i}$ for i<=p1i<=p-1, such $V$ exists since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions (which means that $v_{1},\ldots,v_{p-1}$ are linearly independent).
+Let $V$ be an arbitrary linear transform that maps $e_{i}$ to $v_{i}$ for i<=p1i<=p-1, such $V$ is a group isomorphism between $\mathbb{Z}^{p-1}$ and $V\mathbb{Z}^{p-1}$ since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions (which means that $v_{1},\ldots,v_{p-1}$ are linearly independent).
 Further set
 $A=\{0,v_{1},\ldots,v_{p-1}\},\qquad E=\{0,e_{1},\ldots,e_{p-1}\}.$
-By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the space obtained by applying $V$ to $\operatorname{span}(E)$, which is a sublattice since $\operatorname{span}(E)=\mathbb{Z}^{p-1}$.
+By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the sublattice $V\mathbb{Z}^{p-1}$.
 This implies that $A$ is also tiling in $\mathbb{Z}^{d}$.
@@
 Theorem 2 produces some interesting consequences, for example we can assert that any $3$ points must be tiling as long as they are not on a line.
-On the other hand, this also shows that the condition that $p$ points have to be in general linear positions is tight, since $3$ colinear points indeed need not be tiling, e.g., $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [19] for a characterization of prime power size tiles in $\mathbb{Z}$).
+On the other hand, this also shows that the condition that $p$ points have to be in general linear positions is tight, since $3$ colinear points indeed need not be tiling, e.g., $\{0,3,4\}$ does not tile $\mathbb{Z}$ (see also [19] for a characterization of tiles with prime power sizes in $\mathbb{Z}$).
 Acknowldgement